OpenStudy (anonymous):

Simplify 1/2 (g − 16) A.) g − 8 B.) g + 8 C.) g + 16 D.) g − 16

OpenStudy (anonymous):

if there's \[\LARGE \frac{1}{2}\cdot (g-16)\] none of the multiple choices is correct !

OpenStudy (anonymous):

maybe there's \[\LARGE \frac{1}{2}\cdot (2g-16)\] ??

OpenStudy (anonymous):

no 1/2 (g - 16)

OpenStudy (anonymous):

that would be \[\LARGE \frac12 \cdot (g-16)=\frac{x-16}{2}=\] \[\LARGE \frac x2-\frac{16}{2}=\boxed{\frac x2-8} \]

OpenStudy (anonymous):

sorry is "g" instead of "x"